### v2

# Predicted High-Pressure Hot Superconductivity in Li2CaH16 and Li2CaH17 Phases that Resemble the Type-II Clathrate Structure

Abstract
High-temperature high-pressure superconducting hydrides are typically characterized by cage-like hydrogenic lattices filled with electropositive metal atoms. Here, density functional theory based evolutionary crystal structure searches find two phases that possess these geometric features and are related to the Type-II clathrate structure. In these Li2CaH16 and Li2CaH17 phases the calcium atom occupies the larger cage and the lithium atom the smaller one. The highest superconducting critical temperatures predicted within the isotropic Eliashberg formalism, 330 K at 350 GPa for Li2CaH16 and 370 K at 300 GPa for Li2CaH17, suggest these structures are another example of hot superconducting hydrides. As pressure is lowered the cage-like lattices distort with the emergence of quasimolecular hydrogenic motifs; nonetheless Li2CaH17 is predicted to be superconducting down to 160 GPa at 205 K.
1 Introduction
In 1968, Neil Ashcroft predicted that metallic hydrogen – the hypothetical ground state of the lightest element under pressure – could behave as a high-temperature superconductor 1. The realization of this state has proven elusive, with the most recently estimated pressure needed being over 570 GPa 2. Nearly four decades later Ashcroft described how the pressures required to metallize hydrogen could be decreased via addition of a second element that "chemically precompressed" the system 3. As early as 2015, experimental studies supporting Ashcroft’s prediction came to the fore. These included measurements of high superconducting critical temperatures (s) in phases with unique stoichiometries, stable only under pressure, including H3S ( = 203 K, 150 GPa) 4, LaH10 ( = 260 K, 200 GPa) 5, 6, YH9 ( = 262 K, 182 GPa) 7, YH6 ( = 224 K, 166 GPa) 8, CaH6 ( = 210-215 K, 160-172 GPa) 9, 10, 11, and CeH9 12, 13, 14 ( = 100 K, 166 GPa). Recently, local magnetometry experiments on CeH9, performed via nitrogen-vacancy centers embedded within the diamond anvil, provided evidence of the Meissner effect 15.
The immense pressures required to stabilize the superconducting binary hydrides can only be reached in diamond anvil cells and come coupled with experimental intricacy and expense, making theoretical calculations an appealing method to both determine systems of interest and interpret experimental results 16, 17. Density functional theory (DFT) calculations have been found capable of predicting superconducting properties of conventional superconductors, and crystal structure prediction (CSP) techniques have been successful in uncovering promising synthetic targets (for unit cells that are not too complex) 18. The chemical space has been thoroughly scoured for fruitful high-temperature superconducting binary hydrides 19, 20, 21. In contrast, ternary hydrides present relatively unexplored possibilities, however predictions for this class of hydrides are significantly more challenging 22. The vast chemical space means that an increased number of stoichiometries must be considered for each elemental combination, with a more complex potential energy landscape that requires a greater number of structures be optimized in the CSP runs 23. To overcome these challenges, machine learning potentials have been used to roughly explore the phase diagram prior to performing more costly DFT optimizations 24, and high-throughput calculations have been employed to study elemental substitutions on previously-discovered superconducting templates 25. The introduction of a third element could potentially be used to lower the stabilization pressure, such as in synthesized LaBeH8 ( = 110 K, 80 GPa) 26 and predicted BaSiH8 ( = 71 K, 3 GPa) 25, or increase the superconducting critical temperature as in the synthesized (La,Ce)H9 ( = 148-178 K, 97-172 GPa) 27 and predicted Li2ScH20 ( = 242 K, 300 GPa) 28. The former strategy is usually accomplished by using one electropositive element and one -block element, whereas the latter strategy usually requires the combination of two electropositive elements. Recently, superconductivity in a quaternary hydride, (La,Y,Ce)H10, has been reported with a maximum of 190 K at 112 GPa 29.
The hydrogenic lattices of the binary hydride superconductors of an electropositive element that have very high s generally assume cage-like structures within which a metal atom is encapsulated. Several stoichiometries are well known. The MH6 sodalite-like clathrate, first exemplified by CaH6 9, 10, 11, has a metal atom, M, centered in an H24 cage consisting of six square and eight hexagonal faces. The MH9 structure, which surrounds the metal atom with an H29 cage of six irregular squares, six pentagons and six hexagons, is found in numerous rare-earth compounds including synthesized YH9 7 and CeH9 12, 13, 15. Further increasing the hydrogen content, MH10 stoichiometries, where the metal atom sits within hydrogenic polyhedra of six squares and twelve hexagons, are perhaps the most well-known owing to the synthesis of the LaH10 5, 6 superhydride. The structure of the hydrogen lattice in these MH10 systems is known as “AST” in the clathrate and zeolite community 30. More recently, larger hydride cages have been predicted to encapsulate rare earth metals. These include MH18 (M=Y, La, Ce, Ac, Th) stoichiometries with cubic ( and ) space groups, possessing H36 cages, with a maximum predicted of 329 K at 350 GPa for CeH18 31. Can we draw inspiration from these systems and other occurring clathrates, specifically with phases that contain two electropositive elements?
Looking across the periodic table to other cage-forming elements we encounter silicon, which is able to form large polyhedral structures of sp3 bonded Si atoms that are filled with metal atoms, characterizing the silicon clathrates (SiCL). Two of the most well studied SiCLs are the type-I M8Si24 (the Weaire-Phelan structure), which can be interpreted as two Si20 and 6 Si24 cages filled with a metal atom, and the type-II Mx≤24Si136 structure, which is made of 16 Si20 and 8 Si28 cages with a variable number of metal atoms that can be controlled by the synthesis conditions 32, 33, 34, 35, 36. Several of the known silicon clathrates are superconductors, for example the type-I Ba8Si46 37, with an ambient pressure of 8 K. Recent experiments report the synthesis of (Ba/La/Eu)4H23 38, 39, 40 high pressure hydrides in the type-I structure, with untested s, and Lu4H23 41, with a of 73 K at 218 GPa. Furthermore, predicted Li2(La/Y)H17 28 both match the type-II structure, with the highest being 156 K at 160 GPa for Li2LaH17. These structure types have also been explored for Li2NaH16, Li2NaH17, and LiNa3H23 42, predicting a maximum of 340 K at 300 GPa for Li2NaH17. Prior to these structural correlations being drawn, a computational study predicted that when the MgH16 43 structure was doped with Li, an Li2MgH16 “hot superconductor” 44, which differs from the type-II clathrate structure by a single unoccupied Wyckoff site (a clathrate atom at the (8a) 1/8, 1/8, 1/8 site), possessing a record calculated maximum of 473 K at 250 GPa could be stabilized 44.
Herein, through careful DFT calculations we investigated if a high-temperature superconductor analogous to Li2MgH16 44, but where Mg is replaced by Ca, could be stabilized under pressure. Because the type-I and type-II SiCL structures each possess two cages with distinct sizes, we hypothesized they would be particularly well suited templates for incorporating mixed metal atoms, and in our specific case lithium would fill the smaller cages and calcium the larger ones. Lithium is both electropositive, donating charge to the surrounding hydrogen lattice, and light, allowing for engaging in coupled high-frequency phonons with the hydrogen lattice. Furthermore, as several Ca hydrides are synthesized or predicted superconductors, we intended to improve upon their properties with the addition of Li. With these motivations, we performed CSP searches at megabar pressures for Li2CaHn (). Both stable and metastable phases were discovered, and their electronic structures were analyzed. Two dynamically stable metallic structures, one that is a distortion of the type-II SiCL, and one that can be derived from it, were found, and electron-phonon coupling calculations were performed to estimate their isotropic Eliashberg s: 330 K at 350 GPa for Li2CaH16, and 205 K at 160 GPa or 370 K at 300 GPa for Li2CaH17. This work extends the list of predicted “hot” superconducting ternary hydrides by two.
2 Methods
Exploratory crystal structure prediction (CSP) searches using the open-source XTALOPT 45, 46 evolutionary algorithm (EA) 12th 47 release were performed to identify stable and metastable phases. EA runs were performed for the following stoichiometries: Li2CaHn (=10-20) and Li2Ca at 100, 200, and 300 GPa. The initial generation consisted of random symmetric structures created by the RANDSPG algorithm.48 Duplicate structures were identified via the XTALCOMP algorithm for removal from the breeding pool.49 The breeding operators employed and their relative probabilities were: (i) stripple (50%), (ii) permustrain (35%), and (iii) crossover (15%). The minimum number of generated structures was set to 1500. Each structure optimization consisted of four consecutive geometry optimizations, with each increasing in precision. The final geometry optimization step of the EA search employed a cutoff energy of 400 eV and a -grid generated using the -centered Monkhorst-Pack scheme 50, with the number of divisions along each reciprocal lattice vector chosen such that the product of this number with the real lattice constant was 30 Å. After the EA searches concluded, the three most enthalpically favored unique structures at each stoichiometry and pressure were used for further analysis.
Precise geometry optimizations, electronic structure calculations, electron localization functions (ELF), Bader charges, and first principles molecular dynamics (FPMD) simulations were performed using DFT calculations with the PBE functional as implemented in the Vienna Ab Initio Simulation Package (VASP). 51, 52 We employed a planewave basis set with an energy cutoff of 600 eV, along with PAW potentials 53 where the H 1s1, Li 1s22s12p0 and Ca 3p64s23d0.01 electrons were treated explicitly in all calculations. -meshes were generated using the -centered Monkhorst-Pack scheme 50, and the number of divisions along each reciprocal lattice vector was chosen such that the product of this number with the real lattice constant was 50 Å. All FPMD simulations were performed with a 2 2 2 supercell, a single -point at , an ensemble (constant number of particles, constant volume, constant temperature), Nosé-Hoover thermostat 54, a 0.5 femtosecond timestep, and temperatures corresponding to the predicted s. The crystal orbital Hamilton populations (COHPs) and the negative of the COHP integrated to the Fermi level (-iCOHP) were calculated using the LOBSTER package to analyze bonding. 55, 56
To determine if the predicted structures were dynamically stable, and obtain their zero-point energy, harmonic phonons were calculated using the supercell approach. In the supercell approach, Hellmann-Feynman forces were calculated from a supercell constructed by replicating the optimized structure wherein the atoms had been displaced, and dynamical matrices were computed using the PHONOPY code. 57 Quantum ESPRESSO (QE) was used to obtain the dynamical matrix and electron-phonon coupling (EPC) parameters.58 H, Li, and Ca pseudopotentials obtained from the QE pseudopotential library were generated by the Vanderbilt ultrasoft method with a 1s1 configuration for H, a 1s22s12p0 configuration for Li, and a 3s23p64s23d04p0 configuration for Ca. The PBE functional was employed for all QE calculations. Plane-wave basis set cutoff energies were set to 70 Rydbergs (Ry) for all systems. The Methfessel-Paxton Brillouin-zone smearing of 0.015 Ry was applied.59 The EPC parameter was calculated using Gaussian broadenings in increments of 0.005 Ry ranging from 0.005 to 0.050 Ry. A 15 15 15 -grid with a -mesh of 5 5 5 was employed for Li2CaH17 at 300 GPa. Subsequent calculations on Li2CaH17 utilized a -mesh of 3 3 3. This sparser -mesh resulted in a deviation in of <2%, with a 20 K reduction in , and an overall reduction in computational cost by 80%. This was determined to be sufficiently converged for prediction of EPC properties (Table S7). Therefore, a 15 15 15 -grid with a q-mesh of 3 3 3 was utilized for Li2CaH16. For Li2CaH17 the EPC appeared to be converged with these parameters at a broadening of 0.030 Ry, and Li2CaH16 converged at a broadening of 0.050 Ry.
Critical temperature () values were obtained with a Coulomb repulsion parameter, , of 0.1 by the Allen-Dynes modified McMillan equation for strong coupling,
where is the logarithmic average frequency and the and correction factors for strong coupling and shape dependence are functions of , , and (equaling unity for the weak coupling limit, see Ref. 60). s were also obtained by solving the isotropic Eliashberg equations 61 within the Fermi-surface-restricted (FSR) approximation numerically based on the spectral function, , obtained from the QE calculations. Isotropic solutions of values in hydrides have been found to underestimate s by 10% compared to anisotropic solutions 62, 63, 64, 65, and a recent implementation has made it possible to compare the results of the full bandwidth (FBW) approach to those obtained with the FSR approach for hydrides66. The Matsubara frequency cutoff was set to a value of 20 , the root-mean-square phonon frequency, while the number of frequencies was allowed to vary up to 40,000.
3 Results and Discussion
3.1 Hydrogenic Motifs
Let us take a closer look at the predicted Li2CaHn structures, to better understand the peculiarities of their atomic configurations, and in particular the hydrogenic motifs that are so important for their superconducting properties. Based on the extensive work performed on binary high-pressure hydrides containing electropositive elements 21, 23, it is known that the presence of H- and H anions is detrimental for superconductivity. When these molecules are present, the hydrides are metallized via pressure-induced broadening of the H- and H bonding and anti-bonding bands, resulting in a low density of states (DOS) at the Fermi level () and no or low 67, 68. Typically, compounds with H motifs are predicted to possess intermediate s owing to the metallicity that stems from the partial filling of the antibonding H2 orbitals, and a comparatively high DOS at (). Expanding from molecules into extended hydrogenic lattices, a high can be maintained. Compounds with cage-like lattices that enclathrate electropositive atoms are found to possess large values of , the electron-phonon coupling constant, leading to the highest computed and measured superconducting critical temperatures.69
Analysis of the structures found in our CSP searches reveals these same hydrogenic motifs, and illustrates how they vary as a function of hydrogen content and pressure (Table 1). The Li2CaHn stoichiometries were typically characterized by H- and bent, asymmetric H units, interspersed with H2 molecules at all pressures investigated. The most stable phases featured these same motifs at 100 and 200 GPa, however by 300 GPa incomplete clathrate cages began to emerge, suggesting that both the metal to hydrogen ratio and pressure are key for clathrate formation in the high-pressure hydrides. A clathrate cage is found within Li2CaH16 at 300 GPa, but it dissociates to an incomplete clathrate at 200 GPa and further to a hydrogenic lattice containing discrete H, H2 and H- units by 100 GPa. The global minimum Li2CaH17 phase possesses a clathrate structure at both 300 and 200 GPa, but dissociates to form H2 and H- by 100 GPa. Within Li2CaH18 and Li2CaH19 incomplete clathrates are present at 300 GPa, with H2 units interspersed within the lattice for . At pressures of 100 and 200 GPa, these incomplete clathrates dissociate to form H, H2 and H-. Lastly, Li2CaH20 possesses only H2 units at 100, 200, and 300 GPa. Our observations are in-line with a chemical pressure analysis, which illustrates that clathrate formation in the high-pressure hydrides depends on the size-match between the metal cation and the hydrogenic cluster encapsulating it 70. We postulated that the Li2CaH16 and the Li2CaH17 structures could be candidates for high-temperature, and perhaps “hot” superconductivity, given the proclivity of clathrate-structure-types for large EPC values.
Stoichiometry | 100 GPa | 200 GPa | 300 GPa |
Li2CaH10 | H,H2,H- | H2,H- | H,H2,H- |
Li2CaH11 | H2,H- | H2,H- | H2 |
Li2CaH12 | H2,H- | H2,H- | H,H2,H- |
Li2CaH13 | H2,H- | H2,H- | H2 |
Li2CaH14 | H2,H- | H2,H- | Incomplete Clathrate |
Li2CaH15 | H,H2,H- | H2,H- | Incomplete Clathrate |
Li2CaH16 | H,H2,H- | Incomplete Clathrate | Clathrate |
Li2CaH17 | H2,H- | Clathrate | Clathrate |
Li2CaH18 | H2,H- | H2,H- | Incomplete Clathrate, H2 |
Li2CaH19 | H,H2,H- | H,H2,H- | Incomplete Clathrate |
Li2CaH20 | H2 | H2 | H2 |
3.2 Ternary Phase Diagrams
The most stable structures found in the CSP searches were employed to construct 3D convex hulls (Figures 1, S1-S3) at 100, 200, and 300 GPa at zero-temperature and within the clamped nuclei approximation. These plots are based upon the formation enthalpy as a function of the stoichiometry. By definition the convex hull connects phases whose enthalpies are lower than that of any other phase or linear combination of phases at a particular composition. Thus, the phases that lie on the hull are thermodynamically stable, whereas those that lie above it may be important local minima provided they are close to the hull, are dynamically stable, and are protected by sufficiently high barriers. Therefore, it is often useful to plot the distance of a particular structure from the hull, to determine if it could be an important metastable phase. Because the compositions considered all possessed a 2:1 lithium to calcium ratio while the hydrogen content varied, 2D cuts of the generated convex hulls, corresponding to a plane of increasing H content from Li2Ca at 100 GPa, or 2 Li + Ca at 200 and 300 GPa were also plotted (Figure S5, S6). At 300 GPa a Li2CaH17 phase was predicted to be thermodynamically stable, while the structure closest to, but not on, the hull, Li2CaH16, lay 7 meV/atom above it (Table S2 and Figure 2). Notably, as discussed above, both of these possessed hydrogenic clathrate cages.
The zero-point-energy (ZPE) is often important in determining relative energy orderings for systems containing light elements, with the greatest potential influence on H-bearing phases, as they can have the highest frequency vibrations. For a given pressure the ZPE is affected by symmetry, with lower symmetry systems generally having larger ZPEs, and the nature of the hydrogenic motifs present within the crystal, with dihydrogen molecules having a larger ZPE than hydrogens within extended lattices such as clathrate cages. Indeed, it has been shown that the ZPE can affect the identity of the on-hull structures in high-pressure polyhydrides containing electropositive metal atoms 71. Obtaining the ZPE necessitates performing phonon calculations, which allows for confirmation of dynamic stability, indicated by a lack of imaginary modes.
Phonon calculations, performed in the harmonic approximation, suggested that Li2CaH17, with a hydrogenic clathrate lattice, and Li2CaHn, = (10,12,15,18-20), which did not feature clathrate structures, were dynamically stable at 300 GPa. Li2CaH16, on the other hand, possessed two modes with imaginary frequencies. These imaginary frequencies are related to the acoustic modes, with the majority of the atomic displacement being due to hydrogen. These frequencies were found to become real and positive by 350 GPa (Figure S21), indicating stabilization due to pressure. A first-principles molecular dynamics simulation at 330 K, however, showed that the hydrogenic lattice remained intact, with no molecular H units forming. This suggests one or more of the following: the phase is dynamically stable only at finite temperature; the system may be stabilized by the inclusion of anharmonicity or quantum effects; or the fineness of the grid needed to obtain real phonons is prohibitively computationally expensive. Due to the large size of the unit cell of Li2CaH16, and the associated computational expense, further calculations investigating this discrepancy and determining the lower boundary of dynamic stability were not performed, and the ZPE for this phase was estimated neglecting the imaginary modes. Figure 2(b) illustrates that inclusion of the ZPE lowered the enthalpy of Li2CaH16 so that it joined Li2CaH17 on the 300 GPa hull, but the remaining non-clathrate structures became destabilized, lying at least 40 meV/atom above the hull.
Decreasing pressure moves all of the ternary hydrides progressively further from the hull. Li2CaH16 shifts up to be 16 and 35 meV/atom above it, while Li2CaH17 is calculated to be 9 and 36 meV/atom shy of thermodynamic stability at 200 and 100 GPa, respectively (Figures S1-S6). However, these compounds may still be promising at sub-300 GPa pressures, despite not laying on the hull, provided they are dynamically and thermally stable. Analysis of the zero-temperature DFT-computed enthalpies of the compounds contained in the Materials Project Database led to the estimate that 50% of known inorganic crystalline materials are metastable, with a median metastability energy of 15 meV/atom and a 90 percentile of 67 meV/atom 79. For the specific case of high-pressure hydrides, whose synthesis at high-pressures in diamond anvil cells is often catalyzed by non-equilibrium processes such as thousands-of-Kelvin degree laser heating, numerous examples of synthesized DFT-metastable compounds are known. Some examples include Ca2H577, which lays 20 meV/atom above the hull at 25 GPa, and various PHn phases, which lay 30 meV/atom above the hull within the pressure range 80-225 GPa 80, 81, 82, 83.
3.3 Hydrogen Clathrate Structures
Plots of the enthalpy versus pressure of the low lying structures found in our CSP searches (Figures S7, S8) showed that Li2CaH16 became preferred (over other phases with this stoichiometry) above 270 GPa, while for Li2CaH17 the structure was the most stable above 140 GPa. From the other candidates found in our EA searches, these had the lowest enthalpies up to a maximum investigated pressure of 500 GPa. The propensity for high in compounds with hydrogenic clathrate lattices, the observation that Li2CaH16 and Li2CaH17 both possessed such lattices with a high degree of symmetry, and the finding that both lay on the 300 GPa convex hull upon inclusion of the ZPE, prompted us to examine them in more detail.
Li2CaH16 is isotypic with previously predicted Li2MgH16 and Li2NaH16 42, 44. In it, each Ca atom lies in a cage surrounded by four hexagons and twelve pentagons, made of 28 hydrogen atoms, and each Li atom is enclathrated in an 18 atom hydrogen cage with four hexagons and five pentagons. The size of the cages around each metal atom are commensurate with their radii, with Ca nestled in the larger cage and lithium sitting in the smaller cage, as illustrated in Figure 3. At 350 GPa these polygons are highly regular, with H-H bond lengths of 1.05, 1.08, and 1.12 Å. This symmetry lattice is markedly similar to the type-II SiCL structure, differing by a single Wyckoff position. While the clearest hydrogen motifs in Li2CaH16 are the H18 and H28 cages, we can also view the structure as possessing eight H4 tetrahedra in the conventional cell ( = 8), corresponding to one H4 tetrahedra per formula unit. These H4 tetrahedra are centered around the unfilled Wyckoff position, a clathrate atom at the (8a) 1/8, 1/8, 1/8 site, that if filled would make them isotypic to the type-II SiCL structure.
In Li2CaH17, also shown in Figure 3, each Ca atom is surrounded by four hexagons and twelve pentagons, and each Li atom is once again found in the smaller cage, surrounded by eleven pentagons. The additional hydrogen atom in Li2CaH17 renders the phase nearly isostructural with the type-II SiCLs. In comparison with Li2CaH16, Li2CaH17 possesses H4 tetrahedra “stuffed” with the extra hydrogen atom, which assists in completing the cage around the Li ions. The stuffed hydrogen atom may stabilize this structure by balancing the chemical pressure interactions within the lattice, as has been observed in the H-stuffed H4 tetrahedra of LaH10 70, or by allowing the structure to distort more readily. These distortions can be seen with the reduced symmetry of the faces comprising the polygons that make up the clathrate; each pentagon has H-H bond lengths between 1.08 and 1.14 Å, while each hexagon has H-H bond lengths between 1.02 and 1.14 Å at 300 GPa. Investigation of the Li2CaH17 phase shows that these bonds become less uniform as pressure decreases. The phase was determined to be dynamically stable down to 160 GPa, however the full clathrate structure does not persist to this pressure, due to the H-H bond lengths increasing (e.g. see ELF plot in Figure S20).
Li2CaH17 was not anticipated to be the most stable spacegroup for this stoichiometry. Previous studies have predicted an M3H17 phase, where M is a metal atom, in compounds such as Li2(Sc/La/Y/Na)H17 28, 42. These symmetry M3H17 phases are isostructural with the type-II SiCL, while the Li2CaH17 phase can be seen as a distortion of this structure. The external pressure, internal (or chemical pressure) and the number of valence electrons that are transferred from the metal atoms to the hydrogen cage, and potential metal-hydrogen interactions may be important factors driving this distortion. The external pressure most directly affects the (pressure-volume) term to the enthalpy, often favoring the formation of closed packed structures at higher pressures. The chemical pressure is related to the ionic radius of the metal atom and how it fits into the cage, with ions that are too-small or too-large resulting in distorted lattices 70. The number of valence electrons that are donated from the electropositive element to the hydrogenic lattice can influence the types of hydrogenic motifs that are present 11, and the resulting H-H distances, which are also influenced by interactions such as Ca d H2 backbonding 77. To confirm that the phase is indeed favored over , the cubic structure was built and optimized. Comparison of the 300 GPa enthalpies revealed that the phase possessed a lower internal energy, , but a higher term. The bonds in the phase are more distorted than the highly ordered phase, and these distortions likely contribute to the enthalpic favorability of the structure for the metal atoms studied here. Furthermore, the phase of Li2CaH17 was determined to be both dynamically unstable at 300 GPa, and enthalpically disfavored over the pressure range of 100-500 GPa, so it was therefore not considered for additional calculations.
A similar phenomenon of competing phases has recently been investigated in the LaH10 system 84. While DFT calculations with classical nuclei and neglecting anharmonic effects predicted that an phase becomes unstable with respect to distorted phases including rhombohedral LaH10 below 230 GPa, the inclusion of quantum fluctuations resulted in the dynamic stability of the phase across a wide pressure regime. Moreover, the competing phases all converged to LaH10 when quantum fluctuations were applied in the structural relaxation. While it is possible that quantum fluctuations could stabilize Li2CaH17 over the phase, these calculations were not performed for several reasons. First, an increased measure of non-uniformity among the H-H bonds that make up the clathrate faces is observed for FPMD simulations at 370 K, which is in alignment with the phase, but not the highly symmetric phase. Since the thermal energy provided by the thermostat is far in excess of the stabilization energy due to quantum fluctuations in hydrides, this distortion would be expected if the cubic structure was preferred. Moreover, at no point over the pressure regime did the phase become thermodynamically preferred over the phase, in contrast to the behavior observed for LaH10. Finally, the computational expense required was seen as overly prohibitive, though recently developed machine-learning approaches may make these types of calculations possible in the future 85. Compared to molecular H-H bonds at the same pressure, the H-H bond lengths in both Li2CaH17 and Li2CaH16 are longer (Table 2). Bader analysis showed that the caged Li and Ca atoms donate charge to the nearby hydrogen atoms, which may covalently bond with one another. For Li2CaH17 at 160 GPa, each H atom gains on average 0.16, with Ca losing 1.05 and Li donating a smaller amount of charge, 0.84. When the pressure is increased to 300 GPa, each H atom in Li2CaH17 gains on average 0.15, while Ca loses 0.89 and Li donates 0.80. Thus, increasing pressure appears to even out the amount of charge transferred from both the formally divalent and monovalent metal atoms, perhaps because a shorter distance between the hydrogen atom and calcium can enhance the amount of backdonation between the occupied antibonding hydrogen-lattice states and the Ca d orbitals, as previously found within compressed CaH4 77 and Ca-S-H ternary hydrides 86. The Bader charges calculated for Li2CaH16 at 350 GPa are similar, with H gaining 0.15 on average, and Ca and Li donating 0.90 and 0.78, respectively.
System | Pressure (GPa) | r (Å) | -iCOHP (eV/bond)† | Edges |
Li2CaH16 | 350 | 1.05 | 2.12 (2.04) | 2x |
1.08 | 1.67 (1.98) | 6x , 2x | ||
1.12 | 1.44 (1.60) | 1x | ||
H2 | 350 | 0.75 | 6.37 | |
1.14 | 1.16 | |||
Li2CaH17 | 300 | 1.00 | 2.60 (2.45) | 1x |
1.03 | 2.09 (2.46) | 2x , 1x | ||
1.08 | 2.32 (2.00) | 2x | ||
1.09 | 1.85 (1.83) | 2x | ||
1.13 | 1.23 (1.65) | 6x | ||
1.14 | 1.46 (1.54) | 4x , 2x | ||
H2 | 300 | 0.75 | 6.35 | |
1.14 | 1.13 | |||
CaH6 | 300 | 1.15 | 1.33 | 6x , 4x |
† The -iCOHP values for a hypothetical lattice of hydrogen where no Li or Ca is present, but the H positions are unchanged are provided in parentheses.
To further investigate the bonding within the hydrogenic lattice, the negative of the crystal overlap Hamiltonian population, integrated to the Fermi level (-iCOHP), was calculated for Li2CaH16 and Li2CaH17 at 350 and 300 GPa, respectively, and compared to the -iCOHP for the 2/ phase of molecular hydrogen at the same pressure (Table 2). Unsurprisingly, the intermolecular -iCOHP for 2/ H2 was significantly stronger than that obtained for the shortest distance in the ternary hydrides (6.37 vs. 2.12 eV/bond, and 6.35 vs. 2.60 eV/bond, respectively). Interestingly, though the shortest intramolecular distances within 2/ H2 were nearly identical to the longest H-H bonds in Li2CaH16 and Li2CaH17, the -iCOHPs calculated for these atom pairs in the clathrate cages were somewhat stronger (1.16 vs. 1.44 eV/bond, and 1.13 vs. 1.46 eV/bond, respectively). We can compare the -iCOHPs of the Li2CaH16/17 phases with those computed for the compressed calcium polyhydrides, in particular CaH6, a sodalite-like structure with a hydrogen cage. At 300 GPa, the bond measuring 1.15 Å in CaH6 has an -iCOHP value of 1.33 eV/bond, making it slightly weaker than the 1.14 Å H-H bond in Li2CaH17 at 300 GPa (1.46 eV/bond). Removing the alkali and alkaline earth metal atoms from the lattice, but without any further relaxation, yields a different set of -iCOHP values, illustrating how these electropositive elements affect the bonding in the hydrogen lattice. Though in some cases the interactions are strengthened and in other times they are weakened by the presence of the metal atoms, overall the -iCOHPs computed for the empty cages become more equalized with a somewhat smaller spread.
3.4 Electronic Structure and Superconductivity
Li2CaH16 and Li2CaH17 are calculated to have a high (Figure 4), primarily of hydrogen character, boding well for superconductivity. The contribution of the calcium states to is intermediate, while the contribution of the lithium states is minimal. Therefore, the DOS plots are consistent with a Li+ cation, and a not-fully-ionized calcium, resulting from the back-donation of charge from the hydrogenic lattice to the Ca d orbitals, as supported by the large amount of Ca d states at . In Li2CaH16 at 350 GPa and Li2CaH17 at 300 GPa lies just below a peak in the DOS, suggesting that further electron doping (e.g. by replacing Li via a divalent metal atom with a similar ionic radius, or Ca by a trivalent metal atom) could potentially enhance . Such peaks in the DOS have previously been reported in symmetry Li2NaH17 (just above ) and Li2YH17 (just below ) at 300 GPa 28, 42. Assuming a full transfer of electrons from the electropositive element to the hydrogenic lattice yields three-donated-electrons in Li2NaH17 and four-donated-electrons in Li2CaH17. In both of these systems sits below the peak in the DOS, so more electrons must be added. However, in Li2YH17, with five-donated-electrons, lies just above the peak. Thus, the optimal number of electron required to maximize is between four and five. We find that adding an additional 0.38 per FU (formula unit) (4 /10 FU) to Li2CaH16 at 350 GPa and an additional 0.40 per FU (4 / 10 FU) to Li2CaH17 at 300 GPa would place on that peak. Therefore, a quaternary compound with an approximate Li2Ca0.6M0.4H16/17 stoichiometry, where M is a trivalent metal ion such as Sc, Y, or La may potentially result in a higher than those calculated here, provided the structure is dynamically stable.
The descriptors for high- superconductivity – and the logarithmic average phonon frequency, – were computed for these two phases at pressures where they were dynamically stable (Table 3). The parameter provides the strength of the electron-phonon coupling in conventional superconductors and falls in the exponential in the semi-empirical Allen-Dynes modified McMillan equation often used to estimate (Equation 1). The parameter lies in the pre-factor of Equation 1. A short overview of the various methods that are commonly used to estimate can be found in Ref. 23.
For Li2CaH16, imaginary modes were absent only at 350 GPa, where this phase possessed the largest that we computed. The large of 2.58 necessitated that the Eliashberg equations be solved numerically, and using a value of 0.1 for the Coulomb shielding parameter, , we obtained a of 330 K for Li2CaH16 at this pressure. Though this predicted is over 100 K lower than that computed for the isotypic Li2MgH16 at 250 GPa 44, it surpasses room temperature, rendering Li2CaH16 a “hot” superconductor. Its phonon band structure, phonon projected densities of states and Eliashberg spectral function are shown in Figure 5(a). Because of the mass differences between the three atom types, the phonons can be well separated into regions where the contributions from a particular atom are dominant. Therefore, the total EPC could be partitioned into contributions originating from vibrations in the low-frequency regime, projected mostly onto the calcium atoms (15%), followed by a lithium-dominant region (9%), then by the mainly hydrogenic modes above 750 cm-1 (76%). No high-frequency modes indicative of the presence of molecular hydrogen units were observed, with the highest vibrational frequency being 2385 cm-1, in-line with the previously computed value of 2400 cm-1 for Li2MgH16 at 300 GPa 44. The much higher of Li2MgH16 at 300 GPa compared to that of Li2CaH16 at 350 GPa (3.35 vs. 2.58) is likely in part due to the lower pressure at which it could be stabilized, as was shown to decrease with increasing pressure.44 The larger and lighter mass of Mg as compared to Ca are both factors that result in the higher of Li2MgH16. Another compound that resembled Li2CaH16, LiNa3H23with a type-I SiCL structure, possessed a very similar , and therefore a very similar of 310 K at 350 GPa was predicted 42.
System | Pressure (GPa) | (K) | () | (K) | (K) | ||
Li2CaH16 | 350 | 2.58 | 1123 | 1.23 | 272 | 330 | 0.94 |
Li2CaH17 | 300 | 2.56 | 1314 | 0.99 | 283 | 370 | 1.22 |
Li2CaH17 | 220 | 3.42 | 922 | 0.84 | 286 | 317 | 1.42 |
Li2CaH17 | 160 | 1.72 | 1194 | 0.74 | 179 | 205 | 1.24 |
Let us now turn to Li2CaH17, whose EPC was well above the threshold requiring the use of the Eliashberg formalism in the studied pressure range: 2.56 at 300 GPa and 1.72 at 160 GPa. At 300 GPa the predicted , 370 K, was near the boiling point of water on Earth’s surface, rendering this another example of a “hot” superconductor. This critical temperature, while lower than that predicted for Li2MgH16, has greatly improved upon the of CaH6. At this pressure vibrations that fell in the primarily Ca-based low frequency region contributed 15% to the total EPC, whereas Li-dominant vibrations contributed 4%, and primarily H vibrations contributed 81%. The superconducting properties of Li2CaH17 can be compared directly to Li2ScH17, Li2YH17, Li2LaH17, and Li2NaH17, hydride superconductors with similar structures 28, 42 (Table S8). Li2CaH17 and Li2NaH17 not only stand out as being better metals than Li2ScH17, Li2YH17, and Li2LaH17, but also have much larger EPCs, leading to much higher predicted s. Calculations performed with the Allen-Dynes modified McMillan equation for strong coupling (Equation 1) yielded lower estimates of the .
Both the and the decreased with decreasing pressure for Li2CaH17, although was highest at 220 GPa, where was lowest. Analysis of the structure and the ELF (Figure S19) revealed that the hydrogen clathrate is beginning to break apart into molecular units at this pressure, leading to the large increase in caused by the structural instability. However, these molecular units reduce in-line with previously observed trends 23. The formation of molecular units becomes more pronounced as pressure decreases, and by 160 GPa 50% of the hydrogen atoms constitute H2 or H molecules, with the other half of the hydrogens forming 9-membered rings, decreasing and the calculated even further. The for Li2CaH17 at 160 GPa is calculated to be 205 K, surprisingly similar to that of CaH6 ( = 210-215 K, 160-172 GPa) 9, 10, 11. At 160 GPa, the limit of dynamical stability for Li2CaH17 (Figure S23), its superconducting figure of merit 87, 1.24, is markedly similar to that of H3S at 150 GPa, 1.31, and somewhat larger than that of Li2LaH17 at 160 GPa, 0.95.4, 28 The highest phonon modes for Li2CaH17 are 2275 cm-1 at 300 GPa and 2190 cm-1 at 160 GPa. These modes can be compared to the highest phonon modes for atomic metallic hydrogen at 500 GPa being approximately 2600 cm-1 88, with those of LaH10 being 2300 cm-1 at 300 GPa 5, 6 and for CaH6 being near 2000 cm-1 at 150 GPa11.
4 Conclusions
Previous theoretical calculations have predicted that “hot superconductors”, compounds that are superconducting at room temperature and above, may be stabilized under extreme pressures 42, 31, 44. Interest in these materials stems from the fact that quantum behavior at atmospheric pressures is typically only observed at low temperatures. Herein, crystal structure prediction searches combined with density functional theory calculations predict two new hot superconductors, with critical temperatures, s, as high as 330 K at 350 GPa for Li2CaH16 and 370 K at 300 GPa for Li2CaH17. These ternary hydride phases are predicted to lie on the 300 GPa convex hull (when zero-point-energy effects are included). Li2CaH17 was found to be dynamically stable down to 160 GPa in the harmonic approximation, where its was predicted to be similar to previously-synthesized H3S 4 and CaH69, 10, 11.
The structures of Li2CaH16 and Li2CaH17 can both be described as clathrate-like hydrogenic lattices where the heavier calcium cation occupies the larger H28 cages, and the smaller lithium cation is situated within H18 or H20 encapsulations. Both of these phases are related to the Type-II clathrate structure with Li2CaH17 being a rhombohedral distortion of the symmetry lattices found in the silicon variants, and Li2CaH16 being derived from the ideal Type-II clathrate structure via removal of a single hydrogen atom. We postulate that the in these ternary hydrides could be enhanced by tuning the average valence of the enclathrated electropositive metal atoms so that the Fermi level sits on the peak in the densities of states, resulting in compositions such as Li2Ca0.6M0.4H16/17, where M is a trivalent metal ion.
5 Acknowledgements
Funding for this research was provided by the NSF award DMR-2136038 and the U.S. Department of Energy, National Nuclear Security Administration, through the Chicago-DOE Alliance Center (CDAC) under Cooperative Agreement DE-NA0003975. E.Z. acknowledges the U.S. Department of Energy, Office of Science, Fusion Energy Sciences funding the award entitled High Energy Density Quantum Matter under Award No. DE-SC0020340. Calculations were performed at the Center for Computational Research at SUNY Buffalo (https://hdl.handle.net/10477/79221). M.R. would like to thank Dr. Francesco Belli for discussions regarding quantum fluctuations and electron-phonon coupling.
6 Supporting Information Available
The Supporting Information is available free of charge on the publication website,
http://pubs.acs.org/. It includes the space group at various pressures, distance above the convex hull, relative enthalpies, equation of states, electronic band structures and densities of states, electron localization functions, Bader charges, phonon density of states, computational parameters in EPC calculations, plots of first-principle molecular dynamic simulation results, superconducting properties of selected phases, and structural parameters. The optimized structures were deposited in as cif files in the Theoretical Crystallography Open Database (TCOD IDs: 30000120-30000153).
Notes
The authors declare no competing financial interest.
![[Uncaptioned image]](extracted/5861338/toc_mockup_2.png)